OpenStudy (anonymous):

What number must you add to complete the square? x^2 + 18x = -13

OpenStudy (anonymous):

LInear equation in two variables

OpenStudy (anonymous):

i need the number

OpenStudy (anonymous):

+-9^2

OpenStudy (anonymous):

factorise it firstly x(x - 18) = 0 then if the answer is 0 one of the 2 parts of the equation must equal 0 as anything times 0 equals 0 x * 0 = 0 (* equals times by) therfore x = 0 and (x - 18) = 0 x = 18 hope that helps

OpenStudy (anonymous):

thanks (:

OpenStudy (anonymous):

Your welcome:)

OpenStudy (anonymous):

\[\huge x(x - 18) = 0\] then if the answer is 0 one of the 2 parts of the equation must equal 0 as anything times 0 equals 0 \[\huge x * 0 = 0\] (* equals times by) \[\huge \therefore \quad x = 0 \] and (x - 18) = 0 \[\huge x = 18 \leftarrow ANSWER\]

OpenStudy (anonymous):

@Isaacisbeast123 copied from http://answers.yahoo.com/question/index?qid=20100714075357AAb6Fgo

OpenStudy (anonymous):

@lgbasallote vote me